Find an equation of the line satisfying the given conditions.
Through the orgin with slope 7

Answer by stanbon(75887)

Find an equation of the line satisfying the given conditions.
Through the orgin with slope 7
-----------------
Form of the equation is y = mx + b
You are told that m=7 and y = 0 when x = 0
So solve for "b".
0 = 7*0 + b
0 = 0 + b
b = 0
----------------
Equation:
y = 7x+0
y = 7x
============
Cheers,
Stan H.